# Reflected multi-entropy and its holographic dual

### v1

@@ IV.1.2 Field theoretical calculation
On the other hand, the holographic result is given by minimizing $L[\mathcal{W}](a,b,r)$ in (22). The partial derivatives of $\frac{c}{3}L[\mathcal{W}](a,b,r)$ with respect to $a,b,r$ are numerically poltted in FIG. 6 and compared with the field theoretical results. The two results match precisely.
IV.2 Finite temperature case
@@
In this subsection we consider reflected multi-entropy in CFT at finite temperature, the bulk dual of which is BTZ black hole [46]
@@ IV.2.1 Holographic calculation
IV.2.2 Field theoretical calculation
From the CFT perspective, we need $\mathcal{Z}_{n^{\mathtt{q}-1},m}$, which can be obtained from the six-point function on a cylinder of infinite length with period $\beta_{\text{CFT}}$
$\sigma_{1}\sigma_{2}\sigma_{3}\sigma_{4}\sigma_{5}\sigma_{6}\right>_{\text{cylinder}}\ ,$ (34)
@@ IV.2.2 Field theoretical calculation
which sends $\{w_{1},w_{2},w_{6}\}$ to $\{\infty,0,1\}$ and $\{w_{3},w_{4},w_{5}\}$ to $\{z_{3},z_{4},z_{5}\}$. Following the same method in Appendix B, we can numerically evaluate the derivative of the reflected multi-entropy with respect to $a$ and compare it with the holographic result (33) as plotted in FIG. 8. Again these two results match precisely.
@@
Figure 8: The partial derivatives of $\frac{c}{3}L[\mathcal{W}]$ and $S_{R}^{(\mathtt{q}=3)}$ (divided by $c/3$) with respect to $a$. We take $y_{H}=5$, $b=1000$, $\epsilon=0.02$ and $a$ ranging from $1$ to $2$.
V Discussion
In this letter we introduced reflected multi-entropy as a multipartite measure for mixed state by combining the recent proposal of multi-entropy and canonical purification. This is a new multipartite generalization of reflected entropy, in contrast to earlier attempts [33, 34] which only involve entanglement entropy and canonical purifications (possibly many times). Remarkably, reflected multi-entropy can be computed by replica trick in CFTs at both zero temperature and finite temperature and we find precise match between CFT results and the holographic results. Our results strongly support a new class of duality between multipartite measures for mixed states and spacetime geometry. A couple of future questions are listed in order: First, it is straightforward to generalize our discussion to multiple canonical purifications. Any multi-entropy with different number of parties for the final pure state provides a multipartite measure for the original mixed state. Reflected multi-entropy is singled out for its simplicity. Second, generalize our holographic conjecture to time dependent cases. Since both reflected entropy and multi-entropy are expected to have covariant generalizations [13, 35], we expect that the covariant reflected multi-entropy is obtained as the area of the minimal extremal surface web rather than the global minimal one. Last but not least, it would be interesting to consider the bulk quantum corrections [47, 48] to our entropy to see if there is a sub-leading correction as a bulk reflected multi-entropy.
Acknowledgement. We are grateful for useful discussions with our group members in Fudan University. We would like to thank Jinwei Chu for helpful discussions. This work is supported by NSFC grant 12375063. This work is also sponsored by Natural Science Foundation of Shanghai (21ZR1409800) as well as Shanghai Talent Development Fund.
Appendix A Derivation of conformal weights (25) and (26)
Figure 9: Left: The path integral representation of the reduced density matrix $\rho_{ABC}$. Right: The path integral representation of |$\ket{\rho_{ABC}^{m/2}}$, which is obtained from the canonical purification of $\rho_{ABC}$. Here $m$ should be even and it will be analytically continued to 1 in the end.
Figure 10: The path integral representation of $\mathcal{Z}_{n^{\mathtt{q}-1},m}$. Here we take $\mathtt{q}=3$, $n=3$ and $m=2$. The part inside the black dashed box is the density matrix |$\ket{\rho_{ABC}^{2/2}}\bra{\rho_{ABC}^{2/2}}$ |. In order to make the diagram clear, we omitted some gluing. For example, the black vertical line in the box denotes that the lower edge of subsystem $A$ in the 1st copy is glued to the upper edge of $A$ in the 2nd copy and we omit the line denoting the gluing between the upper edge of $A$ in the 1st copy and the lower edge of $A$ in the 2nd copy. For $m=2$, if the lower edge of a subsystem in the i-th copy is glued to the upper edge of this subsystem in the j-th copy, the reverse is always true.
In this appendix we derive the conformal weights (25) and (26) from the path integral representation of $\mathcal{Z}_{n^{\mathtt{q}-1},m}$. We consider the $\{\mathtt{q}=3,n=3,m=2\}$ case as an example. Replica trick in FIG. 10 is determined by six twist operators $\sigma_{i=1,\dots,6}$ located at the endpoints of the subsystems $A,B,C$ as shown in the top left corner of FIG. 9.
From FIG. 10 we can read
$\begin{split}\Sigma_{A}&=(1,2)(3,4)(5,6)(7,8)(9,10)(11,12)(13,14)(15,16)(17,18)\ ;\\ \Sigma_{B}&=(2,7)(8,13)(14,1)(4,9)(10,15)(16,3)(6,11)(12,17)(18,5)\ ;\\ \Sigma_{C}&=(2,3)(4,5)(6,1)(8,9)(10,11)(12,7)(14,15)(16,17)(18,13)\ .\\ \end{split}$ (37)
Then we have
Therefore, the conformal weights of $\sigma_{i}$ is [^6] The conformal weight of twist operator is given by [49] $h=\frac{c}{24}\sum\nolimits_{k}p_{k}\left(k-\frac{1}{k}\right)\ ,$ with $p_{k}$ the number of $k$-cycles. $h_{i}=(c/24)\left(2-1/2\right)\times 9$, which is consistent with (25). For general $\{\mathtt{q},n,m\}$, since there are only $m$-cycles in $\sigma_{i=1,\dots,6}$, the conformal weight of $h_{i}$ is thus given by (25).
$\begin{split}\sigma_{16}&=\sigma_{1}\sigma_{6}=\Sigma_{A}\Sigma_{C}^{-1}=(1,5,3)(2,4,6)(7,11,9)(8,10,12)(13,17,15)(14,16,18)\ ;\\ \sigma_{23}&=\sigma_{2}\sigma_{3}=\Sigma_{A}^{-1}\Sigma_{B}=(1,13,7)(2,8,14)(3,15,9)(4,10,16)(5,17,11)(6,12,18)\ ;\\ \sigma_{45}&=\sigma_{4}\sigma_{5}=\Sigma_{B}^{-1}\Sigma_{C}=(1,11,15)(2,16,12)(3,7,17)(4,18,8)(5,9,13)(6,14,10)\ .\\ \end{split}$ (39)
Therefore, the conformal weights of $\sigma_{a}$ are $h_{a}=(c/24)\left(3-1/3\right)\times 6$, which is consistent with (26). Now we generalize this result to general $\left\{\mathtt{q},n,m\right\}$ case, where we have $mn^{\mathtt{q-1}}$ copies and they form a $\underbrace{n\times n\times\cdots\times n}_{\mathtt{q}-1}$ lattice with $m$ copies at each cell point. We focus on the row containing the first cell point [^7] This row contains $nm$ replicas in total. For $n=3$ and $m=2$ case, this row corresponds to the green shaded part in FIG. 10. and consider the product of the parts of $\Sigma_{A}^{-1}$ and $\Sigma_{B}$ restricted to the cell points in this row. We use $\tilde{\Sigma}_{A}$ and $\tilde{\Sigma}_{B}$ to denote the parts of $\Sigma_{A}$ and $\Sigma_{B}$ restricted to this row, and they are given by (since $m$ is even, here we take $m=2\ell$)
The product $\tilde{\Sigma}_{A}^{-1}\tilde{\Sigma}_{B}$ is
Appendix B Numerical computation of the six-point function (27)
In this appendix we give a numerical calculation of the six point function (27) with the six points located at $x_{i}$ (19) and the relevant conformal weights given by (25) and (26). Utilizing the conformal transformation
$x_{i}\to z_{i}=\frac{(x_{i}-x_{2})(x_{1}-x_{6})}{(x_{2}-x_{6})(x_{i}-x_{1})}\ ,$ (42)
can be approximated by
Here $f$ is called the semiclassical block, which can be obtained by solving the following monodromy problem. Consider the second-order differential equation
$\psi^{\prime\prime}(z)+T(z)\psi(z)=0\ ,$ (46)
where $T(z)$ is given by
with $c_{i}$ the accessory parameters satisfying
$\begin{pmatrix}\psi_{1}\\ \psi_{2}\end{pmatrix}\rightarrow M\begin{pmatrix}\psi_{1}\\ \psi_{2}\end{pmatrix}\ .$ (49)
Figure 11: Dominant fusion channel of the six-point function (43).
where $M_{a=16}$ denotes the $2\times 2$ monodromy matrix for the cycle $\gamma_{16}$ enclosing $z_{1}$ and $z_{6}$, and similarly for $a=23,45$; $h_{16}$ is the conformal weight of the leading operator in the OPE contraction of $\sigma_{1}(z_{1})$ and $\sigma_{6}(z_{6})$ as shown in FIG. 11, similarly for $a=23,45$. From (25) we know that $h_{16}=h_{23}=h_{45}=\frac{c}{12}\left(n^{2}-1\right)$.
The relation between the semi-classical block $f$ and the accessory parameters $c_{i}$ are
$\partial f/\partial z_{i}=c_{i}\ .$ (51)
@@ Appendix B Numerical computation of the six-point function (27)
$\begin{split}\frac{\partial S_{R}^{(\mathtt{q})}}{\partial\xi}&=\lim_{n\to 1}\frac{1}{n-1}\frac{c}{3}\sum_{i=1}^{6}\frac{\partial f}{\partial z_{i}}\frac{\partial z_{i}}{\partial\xi}\\ &=\lim_{n\to 1}\frac{1}{n-1}\frac{c}{3}\sum_{i=1}^{6}c_{i}\frac{\partial z_{i}}{\partial\xi}\ .\end{split}$ (52)

### v2

@@ Abstract
On the other hand, the holographic result is given by minimizing $L[\mathcal{W}](a,b,r)$ in (19). The partial derivatives of $(c/3)L[\mathcal{W}](a,b,r)$ with respect to $a,b,r$ are plotted in Fig. 5 and compared with the field-theoretical results. These two results match precisely. Notice that we are comparing the derivatives of the entropy. To match the entropy itself, we have to further include the contributions of the operator product expansion coefficients in the CFT calculation [48].
Figure 5: Comparison of the partial derivatives of $\frac{c}{3}L[\mathcal{W}]$ and $S_{R}^{(\mathtt{q}=3)}$ (divided by $c/3$) with respect to $\xi=a,b,r$ at zero temperature. We take $a=20$, $b=100$, and $r$ ranging from 11 to $10$. The holographic results and the field-theoretical results match well (the difference between them is within 0.4%).
We also examine the agreement between the field-theoretical and holographic results at finite temperature. From the CFT perspective, we compute the six-point function on an infinitely long cylinder with period $\beta_{\text{CFT}}$ while on the gravity side, the bulk dual web length is evaluated in the Bañados-Teitelboim-Zanelli (BTZ) black hole background [54] [48]. We find that the reflected multi-entropy decreases with increasing temperature, indicating that thermal effects tend to wash out multipartite correlations. In the context of thermal systems, our proposal provides a concrete multipartite measure, whose potential connections to quantum chaos and information scrambling warrant further investigation. Another promising application of reflected multi-entropy is to explore whether it can serve as a diagnostic of topological order in mixed states [55, 56].
.5 Discussion
In this Letter we introduced reflected multi-entropy as a multipartite measure for mixed state by combining the recent proposal of multi-entropy and canonical purification and found good agreement between the CFT results and the holographic results. One may worry about the potential bulk replica symmetry breaking for multi-entropy. While the bulk replica symmetry was assumed in the original holographic derivation [32, 31], recent work [27] shows this fails for $n>2$  [^7] For tripartite case, [58] presents the following constraint $\frac{1}{m_{AB}}+\frac{1}{m_{BC}}+\frac{1}{m_{CA}}>1\ ,\quad m_{AB,BC,CA}\in\mathbb{Z}_{+}\ ,$ with $m_{AB}$ the order of $\sigma_{A}^{-1}\sigma_{B}$. For the $n$-th Rényi multi-entropy ($m_{AB}=m_{BC}=m_{CA}=n$), this inequality is violated for $n>2$, indicating replica symmetry breaking. . A similar issue arises for reflected multi-entropy with $n>2$. Nevertheless, our holographic calculation focuses on the $n=1$ case, evading the replica symmetry-breaking regime and ensuring the validity of our results. Our findings strongly support a new class of duality between multipartite measures for mixed states and spacetime geometry. Moreover, there are two natural extensions of our work. First, our discussion can be generalized straightforward to multiple canonical purifications. Second, similar to the covariant generalizations of reflected entropy and multi-entropy [13, 31], we expect that the “covariant reflected multi-entropy” is obtained as the area of the minimal extremal surface web rather than the global minimal one.
Note added. Recently, the authors in [59] found a special quantum state where the reflected entropy is not monotonically decreasing under partial trace thus fails to be a correlation measure. However, the reflected entropy remains a valid correlation measure for holographic states, as emphasized in the same paper. See also Ref. [60] for the monotonicity property of reflected entropy in free fields.
Acknowledgements. We are grateful for useful discussions with our group members in Fudan University. We would like to thank Jinwei Chu for helpful discussions. This work is supported by NSFC grant 12375063. This work is also sponsored by Natural Science Foundation of Shanghai (21ZR1409800) as well as Shanghai Talent Development Fund.
I Supplemental Material
I.1 Derivation of conformal weights
In this section we derive the conformal weights of $\sigma_{i=1,\dots,2\mathrm{q}}$ and $\sigma_{a=ij}$
$h_{i=1,\dots,2\mathtt{q}}=\frac{c}{24}\left(m-\frac{1}{m}\right)n^{\mathtt{q}-1}\ ,$ (25)
$h_{a=ij}=\frac{c}{12}\left(n^{\mathtt{q}-1}-n^{\mathtt{q}-3}\right)$ (26)
from the path integral representation of $\mathcal{Z}_{n^{\mathtt{q}-1},m}$. We consider the $\{\mathtt{q}=3,n=3,m=2\}$ case as an example. Replica trick in Fig. 7 is determined by six twist operators $\sigma_{i=1,\dots,6}$ located at the endpoints of the subsystems $A,B,C$ (see top left corner of Fig. 6).
From Fig. 7 we can read
[…]
Then we have
Therefore, the conformal weights of $\sigma_{i}$  [^8] The conformal weight of twist operator is given by [47] $h=\frac{c}{24}\sum\nolimits_{k}p_{k}\left(k-\frac{1}{k}\right)\ ,$ with $p_{k}$ the number of $k$-cycles. is $h_{i}=(c/24)\left(2-1/2\right)\times 9$, which is consistent with (25). For general $\{\mathtt{q},n,m\}$, since there are only $m$-cycles in $\sigma_{i=1,\dots,2\mathrm{q}}$, their conformal weights $h_{i}$ is thus given by (25).
[…]
Figure 7: The path integral representation of $\mathcal{Z}_{n^{\mathtt{q}-1},m}$. Here we take $\mathtt{q}=3$, $n=3$ and $m=2$. The part inside the dashed box is the density matrix $|\ket{\rho_{ABC}^{2/2}}\bra{\rho_{ABC}^{2/2}}$. Some gluings involving replicas have not been shown. In this figure, if the lower edge of a subsystem in the i-th copy is glued to the upper edge in the j-th copy, then the upper edge of the same subsystem in the i-th copy is also glued to the lower edge in the j-th copy.
[…]
Therefore, the conformal weights of $\sigma_{a}$ are $h_{a}=(c/24)\left(3-1/3\right)\times 6$, which are consistent with (26). Now we generalize this result to general $\left\{\mathtt{q},n,m\right\}$ case, where we have $mn^{\mathtt{q-1}}$ copies and they form a $\underbrace{n\times n\times\cdots\times n}_{\mathtt{q}-1}$ lattice with $m$ copies at each cell. We focus on the row containing the first cell  [^9] This row contains $nm$ replicas in total. For $n=3$ and $m=2$ case, this row corresponds to the green shaded part in Fig. 7. and consider the product of the parts of $\Sigma_{A}^{-1}$ and $\Sigma_{B}$ restricted to the cells in this row. We use $\tilde{\Sigma}_{A}$ and $\tilde{\Sigma}_{B}$ to denote the parts of $\Sigma_{A}$ and $\Sigma_{B}$ restricted to this row, and they are given by (since $m$ is even, here we take $m=2\ell$)
The product $\tilde{\Sigma}_{A}^{-1}\tilde{\Sigma}_{B}$ is
I.2 CFT computation of the six-point function
The reflected multi-entropy can be calculated by replica trick
$S_{R}^{(\mathtt{q})}=\lim_{m\to 1}\lim_{n\to 1}\frac{1}{1-n}\frac{1}{n^{\mathtt{q}-2}}\log\frac{\mathcal{Z}_{n^{\mathtt{q}-1},m}}{(\mathcal{Z}_{1,m})^{n^{\mathtt{q}-1}}}\ ,$ (32)
with $\mathcal{Z}_{n^{\mathtt{q}-1},m}$ the partition function on a glued Euclidean spacetime with $mn^{\mathtt{q}-1}$ copies, which can be expressed using the twist operator correlation function. For the $\mathrm{q}=3$ case we are considering, it is given by a six-point function
$\mathcal{Z}_{n^{3-1},m}=\left\langle\sigma_{1}\sigma_{2}\sigma_{3}\sigma_{4}\sigma_{5}\sigma_{6}\right\rangle_{\text{CFT}^{\otimes mn^{3-1}}}\ ,$ (33)
with the six points located at $x_{i}$
$\begin{split}x_{6}&=-x_{1}=b\ ,\\ x_{5}&=-x_{2}=a+r\ ,\\ x_{4}&=-x_{3}=a-r\end{split}$ (34)
and the relevant conformal weights given by (25) and (26). In this section we give a numerical calculation of the six-point function (33). Utilizing the conformal transformation
$x_{i}\to z_{i}=\frac{(x_{i}-x_{2})(x_{1}-x_{6})}{(x_{2}-x_{6})(x_{i}-x_{1})}\ ,$ (35)
can be approximated by
Here $f$ is called the semi-classical block, which can be obtained by solving the following monodromy problem. Consider the second-order differential equation
$\psi^{\prime\prime}(z)+T(z)\psi(z)=0\ ,$ (39)
where $T(z)$ is given by
with $c_{i}$ the accessory parameters satisfying
$\begin{pmatrix}\psi_{1}\\ \psi_{2}\end{pmatrix}\rightarrow M\begin{pmatrix}\psi_{1}\\ \psi_{2}\end{pmatrix}\$. (42)
where $M_{a=16}$ denotes the $2\times 2$ monodromy matrix for the cycle $\gamma_{16}$ enclosing $z_{1}$ and $z_{6}$, and similarly for $a=23,45$;
Figure 8: Dominant fusion channel of the 6-point function (36).
$h_{16}$ is the conformal weight of the leading operator in the OPE contraction of $\sigma_{1}(z_{1})$ and $\sigma_{6}(z_{6})$ as shown in Fig. 8, similarly for $a=23,45$. From (26) we know that $h_{16}=h_{23}=h_{45}=\frac{c}{12}\left(n^{2}-1\right)$.
The relation between the semi-classical block $f$ and the accessory parameters $c_{i}$ is given by
$\partial f/\partial z_{i}=c_{i}\ .$ (44)
$\begin{split}\frac{\partial S_{R}^{(\mathtt{q})}}{\partial\xi}&=\lim_{n\to 1}\frac{1}{n-1}\frac{1}{n^{3-2}}\frac{c}{3}\sum_{i=1}^{6}\frac{\partial f}{\partial z_{i}}\frac{\partial z_{i}}{\partial\xi}\\ &=\lim_{n\to 1}\frac{1}{n-1}\frac{1}{n^{3-2}}\frac{c}{3}\sum_{i=1}^{6}c_{i}\frac{\partial z_{i}}{\partial\xi}\ .\end{split}$ (45)
I.3 Reflected multi-entropy at finite temperature
@@
In this section we consider reflected multi-entropy in CFT at finite temperature, the bulk dual of which is the BTZ black hole [54]
@@
I.3.2 Field-theoretical calculation.
From the CFT perspective, we need $\mathcal{Z}_{n^{\mathtt{q}-1},m}$, which can be obtained from the six-point function on a cylinder of infinite length and period $\beta_{\text{CFT}}$
$\left<\sigma_{1}\sigma_{2}\sigma_{3}\sigma_{4}\sigma_{5}\sigma_{6}\right>_{\text{cylinder}}\ ,$ (52)
@@
which sends $\{w_{1},w_{2},w_{6}\}$ to $\{\infty,0,1\}$ and $\{w_{3},w_{4},w_{5}\}$ to $\{z_{3},z_{4},z_{5}\}$. Following the same method as at zero temperature, we can numerically evaluate the derivative of the reflected multi-entropy with respect to $a$ and compare it with the holographic result (51) as plotted in Fig. 10. Again these two results match precisely.
@@
Figure 10: The partial derivatives of $\frac{c}{3}L[\mathcal{W}]$ and $S_{R}^{(\mathtt{q}=3)}$ (divided by $c/3$) with respect to $a$ at different temperature are represented by lines and dots respectively.
The variation of reflected multi-entropy with respect to $y_{H}=\beta/(2\pi)$ for different interval length $a$ is shown in Fig. 11. As the temperature rises, the reflected multi-entropy decreases, which fits our expectations since in thermal CFT, the reflected entropy has similar behavior in the phase where the entanglement wedge cross section ends on the horizon (see Eq. (24) of [10] for example). As the parameter $a$ (which characterizes the size of region $B$) increases, $S_{R}^{(3)}$ increases, indicating that it is a correlation measure for holographic states.
Figure 11: The variation of reflected multi-entropy $S_{R}^{(\mathtt{q}=3)}$ (divided by $c/3$) with respect to $y_{H}$ for different $a$.
I.4 The OPE coefficient
The monodromy method can only determine the derivatives of the six-point function (45). To obtain the full field-theoretical results, we need to fix the OPE coefficients appearing in (37). There are two kinds of OPE coefficients, $c_{i,j}^{a}$ and $c_{a,b}^{c}$, with $i,j,\dots=1,\dots,6$ labeling the twist operators with conformal weights (25) and $a,b,c=16,23,45$ labeling those with conformal weights (26). The OPE coefficient $c_{i,j}^{a}$ is similar to the OPE coefficient of reflected entropy, while $c_{a,b}^{c}$ is the square of the OPE coefficient of multi-entropy. Let us first focus on $c_{i=2,j=3}^{a=23}$, which corresponds to the fusion $\sigma_{2}\sigma_{3}\to\sigma_{23}$. We consider
$\sigma_{2}=\Sigma_{A}^{-1}\ ,\quad\sigma_{3}=\Sigma_{B}\ ,\quad\sigma_{23}=\Sigma_{A}\Sigma_{B}^{-1}\ .$ (55)
The three point function $\langle\sigma_{2}\sigma_{3}\sigma_{23}\rangle$ computes the partition function of the replica space, a $mn^{\mathtt{q}-1}$-sheet Riemann surface, which in fact factorizes into $n^{\mathtt{q-2}}$ number of $mn$-sheet Riemann surfaces. The $mn$-sheet Riemann surface is nothing but the replica geometry appearing in the replica trick of reflected entropy (see appendix C of Ref. [13]), thus we have
$\langle\sigma_{2}\sigma_{3}\sigma_{23}\rangle=\langle\sigma_{g_{A}^{-1}}\sigma_{g_{B}}\sigma_{g_{A}g_{B}^{-1}}\rangle^{n^{\mathtt{q}-2}}\ ,$ (56)
where $g_{A}$ and $g_{B}$ are the twist operators for reflected entropy. Therefore, $c_{2,3}^{23}$ should be the $n^{\mathtt{q}-2}$-th power of the reflected entropy OPE coefficient $C_{n,m}$,
@@ I.4 The OPE coefficient
$\begin{split}&c_{2,3}^{23}=\left(C_{n,m}\right)^{n^{\mathtt{q}-2}}=(2m)^{-4h_{n,\mathtt{q}}}\ ,\\ &\text{with}\ h_{n,\mathtt{q}}=\frac{c}{24}\left(n^{\mathtt{q}-1}-n^{\mathtt{q}-3}\right)\ ,\end{split}$ (57)
which reduces to the OPE coefficient of reflected entropy when $\mathtt{q}=2$.
The other OPE coefficient, $c_{a=16,b=23}^{c=45}$, is the square of the multi-entropy OPE coefficient $C_{n}$ (since in our case the corresponding conformal weights double), whose $n$-derivative can been fixed from holography [32, 34]
$\partial_{n}\log(c_{a,b}^{c})\big|_{n=1}=-2\partial_{n}\log(C_{n})\big|_{n=1}=c\log\left(\frac{2}{\sqrt{3}}\right)\ .$ (58)
With the result (57) and the assumption (58), the six-point function can be completely fixed as following: in the adjacent limit
$|x_{1}-x_{6}|=|x_{2}-x_{3}|=|x_{4}-x_{5}|=2\epsilon\to 0\ ,$ (59)
the six-point function reduces to a three-point function
$\begin{split}\left\langle\sigma_{1}\sigma_{2}\sigma_{3}\sigma_{4}\sigma_{5}\sigma_{6}\right\rangle&\to c_{1,6}^{16}c_{2,3}^{23}c_{4,5}^{45}\left\langle\sigma_{a=16}\sigma_{b=23}\sigma_{c=45}\right\rangle\\ &=(2m)^{-12h_{n,\mathtt{q}=3}}\left\langle\sigma_{a}\sigma_{b}\sigma_{c}\right\rangle\end{split}$ (60)
with $\sigma_{a=16}$ located at $x_{a}=(x_{1}+x_{6})/2$ and similar for $\sigma_{b}$, $\sigma_{c}$. The three-point function is given by
$\left\langle\sigma_{a}\sigma_{b}\sigma_{c}\right\rangle=\frac{c_{a,b}^{c}}{\left(x_{ab}/(2\epsilon)\right)^{2h_{a}+2h_{b}-2h_{c}}\left(x_{bc}/(2\epsilon)\right)^{2h_{b}+2h_{c}-2h_{a}}\left(x_{ca}/(2\epsilon)\right)^{2h_{c}+2h_{a}-2h_{b}}}=c_{a,b}^{c}\left[\frac{(2\epsilon)^{3}}{x_{ab}x_{bc}x_{ca}}\right]^{\frac{c}{6}(n^{2}-1)}\ ,$ (61)
where $x_{ab}\equiv|x_{a}-x_{b}|$ and we take the cutoff to be $2\epsilon$. With the adjacent limit result as an input, we can further use the monodromy method to determine the value of six-point function in all the parameter region (since the monodromy method determines the derivatives of the six-point function). The consistency between the field-theoretical and holographic results can also be verified in the adjacent limit. The field-theoretical result is given by
$\begin{split}S_{R}^{(3)}[n]&=\frac{1}{1-n}\frac{1}{n}\log\left\langle\sigma_{1}\sigma_{2}\sigma_{3}\sigma_{4}\sigma_{5}\sigma_{6}\right\rangle\\ &=-12\frac{1}{1-n}\frac{1}{n}h_{n,\mathtt{q}=3}\log(2m)-\frac{c}{6}\frac{1}{1-n}\frac{1}{n}\left(n^{2}-1\right)\log\frac{x_{ab}x_{bc}x_{ca}}{(2\epsilon)^{3}}+\frac{1}{1-n}\frac{1}{n}\log c_{a,b}^{c}\\ &=\frac{c}{6}\left(1+\frac{1}{n}\right)\log\frac{x_{ab}x_{bc}x_{ca}}{m^{3}\epsilon^{3}}+\frac{1}{1-n}\frac{1}{n}\log c_{a,b}^{c}\ .\end{split}$ (62)
In the $m,n\to 1$ limit,
$\begin{split}S_{R}^{(3)}&=\frac{c}{3}\log\frac{x_{ab}x_{bc}x_{ca}}{\epsilon^{3}}-\partial_{n}\log(c_{a,b}^{c})\big|_{n=1}\\ &=\frac{c}{3}\log\frac{x_{ab}x_{bc}x_{ca}}{\epsilon^{3}}+c\log\left(\frac{2}{\sqrt{3}}\right)\ ,\end{split}$ (63)
which is the same as $(c/3)L[\mathcal{W}]$ (see Ref. [34] for the calculation of $L[\mathcal{W}]$ in the adjacent limit).